# Density-dependent sodium-storage mechanisms in hard carbon materials

### v1

@@ Abstract
Understanding the sodium-storage mechanism in hard carbon (HC) anodes is crucial for advancing sodium-ion battery (SIB) technology. However, the intrinsic complexity of HC microstructures and their interactions with sodium remains poorly understood. We present a multiscale methodology that integrates grand-canonical Monte Carlo (GCMC) simulations with a machine-learning interatomic potential based on the Gaussian approximation potential (GAP) framework to investigate sodium insertion mechanisms in hard carbons with densities ranging from 0.7 to 1.9 g cm-3. Structural and thermodynamic analyses reveal that increasing carbon density reduces pore size and accessibility, thereby modulating the relative contributions of adsorption, intercalation, and pore filling to the overall storage capacity. Low-density carbons favor pore-filling, achieving extremely high capacities at near-zero voltages, whereas high-density carbons primarily store sodium through adsorption and intercalation, leading to lower but more stable capacities. Intermediate-density carbons (1.31.61.3-1.6 g cm-3) provide the most balanced performance, combining moderate capacity ($\approx 400$ mAh g-1), safe operating voltages, and minimal volume expansion (<10<10%). These findings establish a direct correlation between carbon density and electrochemical behavior, providing atomic-scale insight into how hard carbon morphology governs sodium-storage. The proposed framework offers a rational design principle for optimizing HC-based SIB anodes toward high energy density and long-term cycling stability.
@@ I Introduction
The growing demand for renewable energy storage solutions has exposed the limitations of conventional lithium-ion batteries (LIBs), particularly due to the scarcity of lithium in Earth’s crust [Tarascon2020] and the challenges associated with its recycling. These constraints hinder the large-scale deployment of LIBs, motivating the search for alternative technologies. Sodium-ion batteries have emerged [Chayambuka2020] as a promising candidate in this regard, owing to the natural abundance and low cost of sodium, as well as their favorable cycling performance [Hwang2017]. Such features make sodium-ion batteries (SIBs) particularly suitable for grid-scale [Hirsh2020] and large-scale energy storage applications [Slater2013].
@@
Graphitic carbon, the benchmark anode material for LIBs, demonstrates good compatibility with lithium ions and limited volume changes [Schweidler2018]. However, its performance in SIBs is limited by the larger ionic radius of sodium, which induces significant stress and thermodynamic instability during cycling [Slater2013, Yabuuchi2014]. Among alternative anode materials, hard carbon (HC) has attracted considerable attention due to its intrinsic structural stability and favorable electrochemical properties [Chu2023]. Despite its widespread application, the sodium-storage mechanism in HC remains poorly understood. This complexity arises from the heterogeneous nature of HC, which comprises defects, nanoporous domains, and turbostratic graphene layers [Dou2019]. These structural features strongly influence sodium-ion storage and ultimately determine the electrochemical performance of SIBs.
@@
Sodium storage in HC proceeds through three primary mechanisms – adsorption, intercalation, and pore filling – which manifest as distinct features in the charge voltage profile [Chen2022]. In the high-voltage region (typically above 0.20.30.2–0.3 V vs Na+/Na), sodium ions are predominantly adsorbed at defect sites or surface functional groups within the HC framework [Surta2022]. This adsorption process leads to a rapid decrease in voltage as the limited number of high-energy sites become occupied. As the sodium content increases, the system transitions to the intercalation stage, in which sodium ions insert into the voids between graphene-like layers [Huang2024]. This mechanism contributes to the sloping region of the voltage curve and is highly sensitive to interlayer spacing and the degree of structural disorder. At higher levels of sodium insertion, the pore-filling mechanism dominates. Sodium ions aggregate within larger nanopores and voids, forming clusters that give rise to the characteristic low-voltage plateau near zero voltage [Stevens2000]. This stage accounts for a significant portion of the overall capacity, although it may introduce structural stress and influence long-term stability. Recently, Jain et al. [Jian2025] proposed a new interpretation of the voltage plateau origin by tailoring the pore architecture in HCs. Their study demonstrated a linear correlation between nanopore volume and the experimentally measured plateau capacity, highlighting the critical role of porosity in governing sodium-storage behavior.
Theoretically, early studies focused on adsorption, suggesting that sodium ions are anchored onto the surface or defect sites of HC [Datta2014]. Subsequent investigations indicated that an enlarged interlayer spacing, together with structural defects, can weaken the van der Waals interaction between graphene sheets and thereby facilitate Na intercalation [Tsai2015]. In addition, a filling mechanism at low voltage has been proposed, where sodium ions aggregate into quasi-metallic clusters within nanopores [Youn2021]. However, these models oversimplify the problem by neglecting key structural features of HC, including mass-density variations, pore-size distribution, defect density, and turbostratic graphene domains. As a result, fundamental questions remain unresolved, particularly regarding the precise role of microstructural characteristics in governing adsorption, intercalation, and filling processes, as well as the interplay between these mechanisms.
In this work, we perform grand-canonical Monte Carlo (GCMC) simulations to investigate the sodium-storage mechanisms in HC materials with various structures. The GCMC simulations are coupled with a machine-learned interatomic potential (MLIP) based on the Gaussian approximation potential (GAP) framework. We focus on how the density of HC influences sodium adsorption, intercalation, and pore-filling. In particular, we identify the optimal density that enables favorable electrochemical performance without inducing significant volume changes. Gaining a deeper understanding of the relationship between the structural characteristics of HC and sodium-storage behavior is essential for the rational design of high-performance sodium-ion battery (SIB) anodes. Addressing these knowledge gaps will not only accelerate the development of scalable, post-lithium energy-storage technologies but also provide guidance for future innovations in large-scale, low-cost battery systems.
@@ II Methodology
To overcome the limitations of both density functional theory (DFT) and empirical interatomic potentials, there has been a recent paradigm shift to construct computationally efficient potentials by machine learning (ML) from DFT reference data. These ML methods perform high-dimensional fits to the DFT potential energy surface (PES) using a limited set of carefully selected configurations, and subsequently interpolate energies and forces for unseen structures. Unlike empirical potentials, ML models do not typically assume a fixed functional form; instead, they adapt flexibly to the training data. This reduces bias during construction but requires careful fitting and validation to prevent unphysical extrapolations. Among the different ML interatomic potential (MLIP) frameworks, GAP [Bartok2010] is a kernel-based approach founded on Gaussian process regression. In this framework, the PES is regressed using fitting coefficients obtained during training in combination with kernel functions, which are evaluated on the fly. A GAP prediction is made by comparing the atomic descriptor of a current structure to a subset of structures in the database. Each comparison yields a kernel, a measure of similarity, bounded between 0 (two structures are completely different) and 1 (identical). Different descriptors of the atomic structure can be used to describe the atomic environments. The predicted local energy of an atom $i$ is then expressed for a combination of two-body (2b) and many-body (mb) descriptors as
@@
The choice of descriptors is central to an accurate and data-efficient PES representation. We employ a combination of 2b and mb atomic descriptors, in addition to a tabulated splined “core potential” to describe the interaction between two atoms when they are nearby ($\lesssim 1$ Å). The 2b descriptors, constructed with a 5.5 Å cutoff for C–Na and Na–Na interactions, improve the numerical stability of the GAP. To reproduce the underlying DFT PES with higher fidelity, we include many-body descriptors based on the smooth overlap of atomic positions (SOAP) formalism [bartok_2013]. In this work, we use a more data-efficient implementation of the SOAP descriptor [Caro2019].
@@ II.1 Machine-learning interatomic potentials
$\Delta$-learning. Modeling the potential energy surface of the Na-C system is best done by splitting its representation into two components. A high-quality GAP model for carbon is already available [Muhli2021], and it accurately captures the strong (short-range) covalent interactions and weak (long-range) van der Waals interactions in pure carbon as separate terms in the potential. By comparison, interactions between Na and C are simultaneously short-range and weak. Therefore, rather than constructing a full C-Na binary potential from scratch, we adopt a $\Delta$-learning approach [Fujikake2018]: we fit a machine-learning model only to the energy differences induced by Na insertion into carbon matrices. The intercalation energy is defined as
@@
Training database and model fitting. Initial training data were generated by randomly placing Na atoms (up to 10) in all possible C60 isomers [Sure2017], defected graphite and nanoporous structures. The latter were generated by a melt-graphitization-quench protocol and using the GAP model [Muhli2021]. This approach ensures sampling of all the relevant atomic environments that Na ions might encounter in graphitic carbon. For the extended systems, we used a simulation box of 216 atoms to recreate a vast variety of possible defects in graphite and to obtain bended graphene layers and small nanopores. The whole database was computed at the DFT level of theory with the PBE exchange-correlation functional [Perdew1996] using VASP [Kresse1996, Kresse1999]. Single-point DFT energies were obtained using a cutoff energy of 650 eV and one $k$ point. Four valence electrons were explicitly treated for C (2s2 2p2) and seven for Na (2p6 3s1). All the calculations used exactly the same convergence parameters to avoid introducing additional noise in the data.
@@
To refine the model accuracy, we carried out an iterative-training approach. Starting from an initial GAP, we used short MD runs to generate 100 structures at each iteration: 25 graphite, 25 nanoporous carbon (NP-C) with a density of 1.275 g cm-3, 25 NP-C with a density of 1.645 g cm-3, and 25 amorphous carbon (a-C) with a density of 2.48 g cm-3, with 1-25 sodium insertions per structure. We then ran PBE-DFT single-point calculations and added these structures to the growing training database. This procedure was repeated several times to improve the accuracy of the $\Delta$-GAP model, reaching optimal performance after the tenth cycle, with a root mean square deviation of 5 meV/at (Fig. S1). The GAP is freely available from Zenodo for further reuse [Front2025_CNa]. The composition of the final training dataset, detailing the proportion of each structure type, is summarized in Fig. 1.
All the fits were carried out with the gap_fit program [Klawohn2023], part of the QUIP software package [ref_quippy]. Atomic structure generation, manipulation, visualization and MD simulations were done with the Atomic Simulation Environment [Hjorth2017] (ASE), different in-house codes, ovito [Stukowski2010] and the TurboGAP program [ref_turbogap].
@@ II.2 Hard carbon generation
We used the carbon GAP of Muhli et al. [Muhli2021] which integrate Van der Waals corrections to perform molecular dynamics (MD) using the TurboGAP code [ref_turbogap]. First, the system was equilibrated in a randomized liquid state at 9000 K for 10 ps. Next, the liquid was quenched with a linear temperature profile over 6 ps down to 3500 K. The graphitization process consist of annealing stages at 3500 K for 400 ps. Finally, the HC structures were quenched down to 300 K over 20 ps. Temperature control was achieved using the velocity-rescaling Bussi thermostat [Bussi2007] with a time constant of 100 fs. All MD simulations were performed under periodic boundary conditions with a box containing 1728 atoms, using a Verlet integration time step of 1 fs. For this work, five different densities are considered: 0.7, 1.0, 1.3, 1.6 and 1.9 g cm-3 corresponding to cubic box dimensions of 36.63, 32.52, 29.82, 27.81, 26.28 Å per side respectively. These structures are available on Zenodo [Front2025_structures].
Figure 2: 3D snapshots of the five hard carbon samples of different densities. (a) Ring-size distributions, (b) bond-angle distributions, and (c) pore-size distributions.
@@
Structural Characterization. The morphology of the quenched HC samples reveals that most of the carbon forms curved graphene-like fragments, predominantly with sp2 bonding, as quantified in Table S1. These fragments assemble into three-dimensional networks, where edges containing sp motifs and basal planes can be interlinked via sp3-hybridized carbon atoms. To further analyze the structural topology of the graphitic planes, we computed ring-size and bond-angle distributions [Wang2022]. Here, ring size denotes the number of carbon atoms forming a closed covalent loop with the sp2 network. These quantities serve as sensitive metrics for local order and curvature in disordered carbon networks. The analysis was performed using matscipy [Grigorev2024] for all five HC samples. As shown in Fig. 2, the graphene-like layers in all samples are primarily composed of 5-, 6-, and 7-membered carbon rings. With the exception of the 1.9 g cm-3 sample, all densities display similar ring-size distributions characterized by a dominant peak at 6-membered rings, indicative of well-ordered graphitic domains, along with approximately equal populations of 5- and 7-membered rings. In contrast, the 1.9 g cm-3 sample exhibits a higher population of 5-membered rings as compared to 7-membered ones. Complementary information is provided by the bond-angle distributions. All samples feature a sharp peak at 120, a hallmark of regular hexagonal carbon rings in graphitic structures. However, the 1.9 g cm-3 sample shows a less pronounced 120 peak, along with an inflection near 108, which corresponds to pentagonal motifs. This feature is reminiscent of the curvature associated with Buckminsterfullerene-like structures.
@@
Pore-size distributions. Porosity is one of the most critical properties of HC structures, as it directly influences the ability of an electrode material for ion intercalation. Therefore, characterizing porosity and pore structures is a central focus of this work. We calculated the pore-size distributions using the zeo++ code [Pinheiro2013]. A three-dimensional Voronoi network [Willems2012] was constructed to identify the void space within the structures, based on Voronoi decomposition as implemented in the Voro++ library [Rycroft2009]. Monte Carlo sampling was then performed to estimate the accessible pore volume and to compute the largest included sphere radius, defined as the radius of the largest sphere that can be inserted at a given point without overlapping with any of the carbon atoms in the structure. The pore-size distribution (PSD) is shown in Fig. 2c, the PSD histograms display unimodal distributions that resemble normal distributions. With increasing density, the PSDs shift toward smaller pore sizes, with peaks located at 21.5, 11.6, 7.6, 6.0, and 3.4 Å for 0.7, 1.0, 1.3, 1.6, and 1.9 g cm-3, respectively. Note that the value for the 1.9 g cm-3 sample is more indicative of the interlayer spacing in the material than of the porosity. As the density increases, the PSDs also become sharper, indicating that individual pores exhibit less size variation. A more open pore structure corresponds to larger interplanar spacing between graphitic sheets, and thus the presence of larger pores. Consequently, low-density HC may accommodate a larger number of sodium ions than high-density HC.

### v2

@@ Abstract
Understanding the sodium-storage mechanism in hard carbon (HC) anodes is crucial for advancing sodium-ion battery (SIB) technology. However, the intrinsic complexity of HC microstructures and their interactions with sodium remain not fully elucidated. We present a multiscale methodology that integrates grand-canonical Monte Carlo (GCMC) simulations with a machine-learning interatomic potential based on the Gaussian approximation potential (GAP) framework to investigate sodium insertion mechanisms in hard carbons with different levels of porosity, achieved by simulating structural models with densities ranging from 0.7 to 1.9 g cm-3. Structural and thermodynamic analyses reveal the interplay between pore size and accessibility and the relative contributions of adsorption, intercalation, and pore filling to the overall storage capacity. Low-density carbons favor pore-filling, achieving extremely high capacities at near-zero voltages, whereas high-density carbons primarily store sodium through adsorption and intercalation, leading to lower but more stable capacities. Intermediate-density carbons (1.31.61.3-1.6 g cm-3) provide the most balanced performance, combining moderate capacity (480 and 310 mAh g-1), safe operating voltages, and minimal volume expansion (<10<10%). These findings establish a direct correlation between carbon density and electrochemical behavior, providing atomic-scale insight into how hard carbon morphology governs sodium-storage. The proposed framework offers a rational design principle for optimizing HC-based SIB anodes toward high energy density and long-term cycling stability.
@@ I Introduction
The growing demand for renewable energy-storage solutions has exposed the limitations of conventional lithium-ion batteries (LIBs), particularly due to the scarcity of lithium in Earth’s crust [44] and the challenges associated with its recycling. These constraints hinder the large-scale deployment of LIBs, motivating the search for alternative technologies. Sodium-ion batteries have emerged [8] as a promising candidate in this regard, owing to the natural abundance and low cost of sodium, as well as their favorable cycling performance [23]. Such features make sodium-ion batteries (SIBs) particularly suitable for grid-scale [19] and large-scale energy-storage applications [38].
@@
Graphitic carbon, the benchmark anode material for LIBs, demonstrates good compatibility with lithium ions and limited volume changes [35]. However, its performance in SIBs is limited by the larger ionic radius of sodium, which induces significant stress and thermodynamic instability during cycling [38, 50]. Among alternative anode materials, hard carbon (HC) has attracted considerable attention due to its intrinsic structural stability and favorable electrochemical properties [10]. Despite its widespread application, the sodium-storage mechanism in HC remains not fully understood. This complexity arises from the heterogeneous nature of HC, which comprises defects, nanoporous domains, and turbostratic graphene layers [13]. These structural features strongly influence sodium-ion storage and ultimately determine the electrochemical performance of SIBs.
@@
Sodium storage in HC proceeds through three primary mechanisms – adsorption, intercalation, and pore filling – which manifest as distinct features in the charge voltage profile [9]. In the high-voltage region (typically above 0.20.30.2–0.3 V vs Na+/Na), sodium ions are predominantly adsorbed at defect sites or surface functional groups within the HC framework [43]. This adsorption process leads to a rapid decrease in voltage as the limited number of high-energy sites become occupied. As the sodium content increases, the system transitions to the intercalation stage, in which sodium ions insert into the voids between graphene-like layers [21]. This mechanism contributes to the sloping region of the voltage curve and is highly sensitive to interlayer spacing and the degree of structural disorder. At higher levels of sodium insertion, the pore-filling mechanism dominates. Sodium ions aggregate within larger nanopores and voids, forming clusters that give rise to the characteristic low-voltage plateau near zero voltage [39]. This stage accounts for a significant portion of the overall capacity, although it may introduce structural stress and influence long-term stability. Recently, Jian et al. [24] proposed a new interpretation of the voltage plateau origin by tailoring the pore architecture in HCs. Their study demonstrated a linear correlation between nanopore volume and the experimentally measured plateau capacity, highlighting the critical role of porosity in governing sodium-storage behavior.
Early theoretical studies of sodium storage in hard carbon primarily emphasized adsorption mechanisms, proposing that sodium ions preferentially bind to surface sites or structural defects [11]. Subsequent investigations highlighted the role of enlarged interlayer spacing and structural disorder in facilitating sodium intercalation by weakening interlayer van der Waals interactions [45]. At low potentials, a pore-filling mechanism has been proposed, in which sodium ions aggregate into quasi-metallic clusters within nanopores [51]. These studies form the basis of the widely accepted three-stage sodium-storage model in HC. However, most computational investigations of these mechanisms have relied on empirical interatomic potentials [43, 28], which provide a limited and often inaccurate description of Na–C interactions. The limited fidelity of such potentials hampers their ability to capture the diverse chemical environments present in hard carbon [6], particularly under conditions of high sodium loading. Furthermore, these models typically employ simplified structural representations that neglect essential microstructural features, including variations in mass density, pore-size distribution, defect concentration, and the presence of turbostratic graphene domains. As a result, fundamental questions remain unresolved, particularly regarding the precise role of microstructural characteristics in governing adsorption, intercalation, and filling processes, as well as the interplay between these mechanisms. Addressing these limitations requires simulation methodologies capable of accurately describing Na–C interactions while explicitly accounting for the inherent structural heterogeneity of HC.
In this work, we develop a machine-learning–based computational framework with density functional theory (DFT) accuracy to investigate sodium storage in hard carbon. Grand-canonical Monte Carlo (GCMC) simulations are performed using a machine-learning interatomic potential constructed within the Gaussian approximation potential (GAP) framework. This approach enables an accurate and efficient description of Na–C interactions across heterogeneous structural environments. Using this framework, we systematically examine the influence of HC porosity on sodium adsorption, intercalation, and pore filling. In particular, we identify density regimes that enable favorable sodium storage while limiting structural deformation and volume expansion. By establishing quantitative links between hard-carbon microstructure and sodium-storage behavior, this work provides methodological insights relevant to the rational design of HC anodes for sodium-ion batteries and, more broadly, to the modeling of ion storage in disordered electrode materials.
@@ II Methodology
To overcome the limitations of both density functional theory (DFT) and empirical interatomic potentials, there has been a recent paradigm shift to construct computationally efficient potentials by machine learning (ML) from DFT reference data. These ML methods perform high-dimensional fits to the DFT potential energy surface (PES) using a limited set of carefully selected configurations, and subsequently interpolate energies and forces for unseen structures. Unlike empirical potentials, ML models do not typically assume a fixed functional form; instead, they adapt flexibly to the training data. This reduces bias during construction but requires careful fitting and validation to prevent unphysical extrapolations. Among the different ML interatomic potential (MLIP) frameworks, GAP [3] is a kernel-based approach founded on Gaussian process regression. In this framework, the PES is regressed using fitting coefficients obtained during training in combination with kernel functions, which are evaluated on the fly. A GAP prediction is made by comparing the atomic descriptor of a current structure to a subset of structures in the database. Each comparison yields a kernel, a measure of similarity, bounded between 0 (two structures are completely different) and 1 (identical). Different descriptors of the atomic structure can be used to describe the atomic environments. The predicted local energy of an atom $i$ is then expressed for a combination of two-body (2b) and many-body (mb) descriptors as
@@
The choice of descriptors is central to an accurate and data-efficient PES representation. We employ a combination of 2b and mb atomic descriptors, in addition to a tabulated splined “core potential” to describe the interaction between two atoms when they are nearby ($\lesssim 1$ Å). The 2b descriptors, constructed with a 5.5 Å cutoff for C–Na and Na–Na interactions, improve the numerical stability of the GAP. To reproduce the underlying DFT PES with higher fidelity, we include many-body descriptors based on the smooth overlap of atomic positions (SOAP) formalism [2]. In this work, we use a more data-efficient implementation of the SOAP descriptor [7].
@@ II.1 Machine-learning interatomic potentials
$\Delta$-learning. Modeling the potential energy surface of the Na-C system is best done by splitting its representation into two components. A high-quality GAP model for carbon is already available [30], and it accurately captures the strong (short-range) covalent interactions and weak (long-range) van der Waals interactions in pure carbon as separate terms in the potential. By comparison, interactions between Na and C are simultaneously short-range and weak. Therefore, rather than constructing a full C-Na binary potential from scratch, we adopt a $\Delta$-learning approach [17]: we fit a machine-learning model only to the energy differences induced by Na insertion into carbon matrices. The intercalation energy is defined as
@@
Training database and model fitting. Initial training data were generated by randomly placing Na atoms (up to 10) in all possible C60 isomers [42], defected graphite and nanoporous structures. The latter were generated by a melt-graphitization-quench protocol and using the GAP model [30]. This approach ensures sampling of all the relevant atomic environments that Na ions might encounter in graphitic carbon. For the extended systems, we used a simulation box of 216 atoms to recreate a vast variety of possible defects in graphite and to obtain bended graphene layers and small nanopores. The whole database was computed at the DFT level of theory with the PBE exchange-correlation functional [31] using VASP [26, 27]. Single-point DFT energies were obtained using a cutoff energy of 650 eV and one $k$ point. Four valence electrons were explicitly treated for C (2s2 2p2) and seven for Na (2p6 3s1). All the calculations used exactly the same convergence parameters to avoid introducing additional noise in the data.
@@
To refine the model accuracy, we carried out an iterative-training approach. Starting from an initial GAP, we used short MD runs to generate 100 structures at each iteration: 25 graphite, 25 nanoporous carbon (NP-C) with a density of 1.275 g cm-3, 25 NP-C with a density of 1.645 g cm-3, and 25 amorphous carbon (a-C) with a density of 2.48 g cm-3, with 1-25 sodium insertions per structure. We then ran PBE-DFT single-point calculations and added these structures to the growing training database. This procedure was repeated several times to improve the accuracy of the $\Delta$-GAP model, reaching optimal performance after the tenth cycle, with a root mean square deviation of 5 meV/at (Fig. S1). The GAP is freely available from Zenodo for further reuse [15]. The composition of the final training dataset, detailing the proportion of each structure type, is summarized in Fig. 1.
All the fits were carried out with the gap_fit program [25], part of the QUIP software package [33]. Atomic structure generation, manipulation, visualization and MD simulations were done with the Atomic Simulation Environment [20] (ASE), different in-house codes, ovito [41] and the TurboGAP program [46].
@@ II.2 Hard carbon generation
Figure 2: 3D snapshots of five hard carbon samples of different densities. (a) Diffraction patterns, (b) reduced partial distribution functions $G(r)$, and (c) pore-size populations. Standard deviation are represented with horizontal bars. Experimental results are extracted from [37, 36] for the HTW2500 sample (glassy) and from [22] for a HC carbon sample (porous).
@@
We used the carbon GAP of Muhli et al. [30] which integrates Van der Waals corrections to perform molecular dynamics (MD) using the TurboGAP code [46]. First, the system was equilibrated in a randomized liquid state at 9000 K for 10 ps. Next, the liquid was quenched with a linear temperature profile over 6 ps down to 3500 K. The graphitization process consist of annealing stages at 3500 K for 400 ps. Finally, the HC structures were quenched down to 300 K over 20 ps. Temperature control was achieved using the velocity-rescaling Bussi thermostat [5] with a time constant of 100 fs. All MD simulations were performed under periodic boundary conditions with a box containing 1728 atoms, using a Verlet integration time step of 1 fs. To explore the effect of density on structural and electrochemical properties, we generated five target densities: 0.7, 1.0, 1.3, 1.6 and 1.9 g cm-3 corresponding to cubic box lengths of 36.63, 32.52, 29.82, 27.81, 26.28 Å respectively. Recognizing the stochastic nature of the melt–quench protocol and the inherent heterogeneity of HC, we generated ten independent structures per density using different random seeds. Structural properties reported in this work correspond to the mean values across these ensembles, with standard deviations capturing variability. These structures are available on Zenodo [16].
@@
Structural Characterization. The morphology of the quenched HC samples reveals that most of the carbon forms curved graphene-like fragments, predominantly with sp2 bonding, as quantified in Table S1. These fragments assemble into three-dimensional networks, where edges containing sp motifs and basal planes can be interlinked via sp3-hybridized carbon atoms. To further analyze the structural topology of the graphitic planes, we computed ring-size and bond-angle distributions [48]. Here, ring size denotes the number of carbon atoms forming a closed covalent loop with the sp2 network. These quantities serve as sensitive metrics for local order and curvature in disordered carbon networks. The analysis was performed using matscipy [18] for all five HC densities. As shown in Fig. S2(a), the graphene-like layers in all samples are primarily composed of 5-, 6-, and 7-membered carbon rings. All densities display similar ring-size distributions characterized by a dominant peak at 6-membered rings, indicative of well-ordered graphitic domains, along with approximately equal populations of 5- and 7-membered rings. We also note that the small standard deviation confirms the robustness of the simulation protocol. Complementary information is provided by the bond-angle distributions. All samples feature a sharp peak at 120, a hallmark of regular hexagonal carbon rings in graphitic structures along with an inflection near 108. These pentagonal motifs, as heptagon, play an important role in curving graphene fragments in HC and glassy carbons [29]. Neither the ring-size nor the bond-angle distributions show a significant dependence on density. To compare the simulated HC structures with experimental data, we computed diffraction patterns (Fig 2(a)) and reduced pair distribution functions (PDF) (Fig. 2(b)), which serve as metrics for assessing structural similarity. We chose experimental glassy sample HTW2500 reported in ref [36] and a HC carbon sample reported in ref [22] as references. The reciprocal-space diffraction scattering were calculated using the Debye scattering equation as implemented in the Debyer code [12]. The reduced PDF commonly used in experiment is related to the PDF $g(r)$ most often reported in simulation work by $G(r)=4\pi\rho(g(r)-1)$, where $\rho$ is the atomic density. The HTW2500 sample exhibits three main diffraction peaks at $2\theta\sim$ 24, 44 and 80, corresponding to the graphite $\{002\}$, $\{100\}$, and $\{110\}$ reflections. For the lower-density model structures, the $\{002\}$ reflection is absent, owing to reduced and irregular layer stacking resulting from a more open and diffuse arrangement of graphene sheets. The interlayer spacing can be estimated using Bragg’s law. The systematic shift toward smaller scattering angles in Fig. 2(a) with decreasing density clearly indicates an increase in interlayer spacing. Among the simulated structures, only the HC sample with a density of 1.9 g cm-3 exhibits a well-defined $\{002\}$ reflection, although the onset of this peak is already apparent for the 1.6 g cm-3 model. From the position of this reflection, an average interlayer spacing of approximately 3.8 Å can be inferred. In contrast, no significant shift is observed in the in-plane $\{100\}$ reflection, indicating an in-plane lattice parameter $a$ of around 2.43 Å.
@@
Pore-size distributions. Porosity is one of the most critical properties of HC structures, as it directly influences the ability of an electrode material for ion intercalation. Therefore, characterizing porosity is a central focus of this work. We calculated the pore-size distributions using the zeo++ code [32]. A three-dimensional Voronoi network [49] was constructed to identify the void space within the structures, based on Voronoi decomposition as implemented in the Voro++ library [34]. Monte Carlo sampling was then performed to estimate the accessible pore volume and to compute the largest included sphere radius, defined as the radius of the largest sphere that can be inserted at a given point without overlapping with any of the carbon atoms in the structure. We take a probe radius of 1.16 Å (ionic radius of Na+) and 130 000 points for Monte Carlo sampling. From the resulting pore-size distributions, we extracted the mean pore radius for each density. For densities below 1.6 g cm-3, all samples exhibit a trimodal or bimodal pore-size distribution. In contrast, at a density of 1.9 g cm-3, only one third of the samples display a bimodal distribution, while the remaining samples show no accessible pores. Figure S2(a) presents the pore-size distributions for each structure used in the sodium insertion simulations. Porosity is defined as pore size $D$ and pore population as shown in Fig. 2(c). With decreasing density, $D$ shift toward larger pore sizes, where for each density, the larger one exhibit the largest pore population. As the density increases, the PSDs also become sharper and standard deviation smaller, indicating that individual pores exhibit less size variation. A more open pore structure corresponds to larger interplanar spacing between graphitic sheets, and thus the presence of larger pores. Consequently, low-density HC may accommodate a larger number of sodium ions than high-density HC. We note that these PSDs do not explicitly account for percolation or pore connectivity, and some voids may be isolated in individual structures. Consequently, the distributions should be interpreted as indicative of typical structural trends with density rather than as exact descriptors of sodium-accessible pores.